# Condensation of Composite Bosonic Trions in Interacting Bose-Fermi Mixtures

### v3

@@ Abstract
We reveal a quantum coherent state characterized by composite bosonic trions, wherein paired fermions further bind with bosons, in one-dimensional Bose-Fermi mixtures. This phase emerges in two distinct models, featuring onsite boson–fermion attraction that induces negative binding energy for the composite trions. The first is the Bose-Fermi pair–hopping model, in which spinless fermions undergo pair hopping to form preformed pairs. The second is the extended Bose-Fermi Hubbard model, which describes dipolar bosons and fermions with density–density interactions. Notably, the formation of composite trions is independent of the specific pairing interactions of fermions. Extensive density matrix renormalization group calculations demonstrate the quasi-condensation of the composite trions, evidenced by algebraically decaying correlations of trions, gapped single-particle excitations, and suppressed fermion pair correlations. Our findings provide novel insights into a general three-body pairing mechanism among particles with distinct quantum statistics.
@@
Introduction.— Interactions between bosons and fermions are crucial in fundamental processes, spanning diverse fields such as condensed matter physics, quantum optics, and nuclear and particle physics. For instance, protons and neutrons can form nuclei through the absorption or emission of mesons Yukawa (1935). Moreover, the electron-phonon coupling is indispensable in the formation of Cooper pairs, which is crucial for both conventional and unconventional superconductivity Schrieffer (1983); Chen et al. (2021); He et al. (2018); Chen et al. (2014). Remarkably, at room temperature, Bose-Einstein condensates of exciton-polaritons have been realized in semiconductor and organic microcavities Byrnes et al. (2014); Keeling and Kéna-Cohen (2020); Ghosh et al. (2022). Polaritons are hybrid bosonic quasiparticles representing a superposition of an exciton (bound electron-hole pairs) and a cavity photon. Due to their finite lifetime, polaritons inherently possess a non-equilibrium nature. This naturally prompts a fundamental inquiry: Is there a comprehensive theoretical framework elucidating how a fermion pair and a boson can pair to form a composite bosonic trion, yielding a coherent and stable ground state in equilibrium? Furthermore, can this novel phase of matter be realized through quantum simulation?
@@
Over the past two decades, ultracold bosonic and fermionic atoms have proven to be exceptionally versatile platforms for simulating quantum physics Bloch et al. (2008); Gross and Bloch (2017); Schfer et al. (2020). In addition to bosonic trimers associated with well-known Efimov physics Efimov (1973); Kolganova et al. (2017); Kunitski et al. (2015), the formation of fermionic trimers has aroused considerable interest in interacting ultracold Fermi gases Williams et al. (2009); Nakajima et al. (2011); Shi et al. (2014); Orso et al. (2010). Since atomic mixtures are highly controllable, rich phenomena have also been observed in Bose-Fermi mixtures (BFM), including polarons Scazza et al. (2022); Baroni et al. (2024); Fritsche et al. (2021), heteronuclear molecules Stan et al. (2004); Milczewski and Duda (2023); Duda et al. (2023), and double superfluids Ferrier-Barbut et al. (2014); Roy et al. (2017); Yao et al. (2016). Beyond extensive investigations of composite fermions (a fermion plus a boson) Stan et al. (2004); Milczewski and Duda (2023); Duda et al. (2023); Pollet et al. (2006); Cazalilla and Ho (2003); Mathey et al. (2004); Barillier-Pertuisel et al. (2008); Mering and Fleischhauer (2008); Pollet et al. (2008), previous theoretical work has analyzed the formation of trios, comprising a composite fermion and an elementary boson, under three-particle interaction using a diagrammatic formalism Kagan et al. (2004). Recently, an ultracold gas of ${}^{23}Na^{40}K_{2}$ triatomic molecules has been successfully created through an atom–diatomic-molecule Feshbach resonance between fermionic ${}^{23}Na^{40}K$ molecules and fermionic ${}^{40}K$ atoms Yang et al. (2022a, b); Cao et al. (2024). Comparable to ultracold atoms, solid-state BFMs, such as two-dimensional transition metal dichalcogenides, have emerged as attractive platforms for exploring fermionic trions, where excitons bind with free electrons or holes via electrically tunable Feshbach resonances Schwartz et al. (2021); Wagner et al. (2025). These advances in quantum simulation underscore the potential for realizing an exotic pairing process, wherein fermion pairs and bosons bind to form novel composite bosonic trions that can condense. Nevertheless, a comprehensive investigation into the underlying pairing mechanism for composite bosonic trions is still lacking.
@@
In this paper, we investigate the formation of composite bosonic trions, wherein a fermion pair further associates with a boson. We first explore the composite bosonic trion phase within the Bose-Fermi pair–hopping model, where kinetic pair tunneling of spinless fermions Ruhman and Altman (2017); Gotta et al. (2021a, 2022a) induces performed fermion pairs. Numerical evidence reveals that, under onsite boson–fermion attraction, only the trion correlation function displays algebraic decay, indicative of a single gapless bosonic mode with a central charge $c=1$. The quasi-condensation momentum of the composite trions is dictated by the sign of the fermionic pair-hopping amplitude. Furthermore, the stability of the composite trions is reinforced by negative binding energy, gapped single-particle excitations, and suppressed correlations among paired fermions. Our calculations also show that, as the boson–fermion interaction weakens, the composite trion phase transitions smoothly into a two-component Luttinger liquid (TLL), where fermion pairs and bosons each contribute a gapless excitation mode, resulting in an overall central charge of $c=2$.
@@
Building on insights from the pair–hopping model, we generalize our analysis to the extended Bose-Fermi Hubbard model Baranov et al. (2012); Wang et al. (2010); Orignac et al. (2010); Gómez-Lozada et al. (2025), which describes dipolar bosons and fermions interacting via both short-range onsite attractions and long-range density–density interactions (including first- and second–neighbor couplings). Within this framework, we identify a similar composite trion phase that occupies a substantial portion of the phase diagram. Remarkably, the composite trion formation does not depend on the particular physical origin underlying fermion pairing, emphasizing the universal nature of the trion formation mechanism. Our work establishes a versatile platform for realizing composite bosonic trions and reveals a novel three-body pairing mechanism in systems with mixed quantum statistics, paving the way for future explorations of exotic multicomponent quantum phases.
@@
Bose–Fermi pair–hopping model.— The first model we consider incorporates the typical on-site boson-fermion interaction of strength $U_{bf}$ Pollet et al. (2006); Cazalilla and Ho (2003); Mathey et al. (2004); Barillier-Pertuisel et al. (2008); Mering and Fleischhauer (2008); Pollet et al. (2008) and a kinetic component for fermions that exhibits competition between single and paired configurations, as introduced in Ref. Gotta et al. (2021a, b, 2022b); Bariev (1991). The Hamiltonian in a chain of size $L$ is given by
$\begin{split}H=&-\sum_{i}(t_{f}c^{\dagger}_{i}c_{i+1}+t_{b}b^{\dagger}_{i}b_{i+1}+H.c.)+U_{bf}\sum_{i}n^{b}_{i}n^{f}_{i}\\ &-t_{ff}\sum_{i}(c^{\dagger}_{i+1}c^{\dagger}_{i}c_{i}c_{i-1}+H.c.)+\frac{U_{bb}}{2}\sum_{i}n^{b}_{i}(n^{b}_{i}-1)\end{split}$ (1)
@@
where $c^{\dagger}_{i}$ ($b^{\dagger}_{i}$) creates a spinless fermion (a boson) at site $i$, $t$ and $t_{ff}$ denote the tunneling rates for single particles and paired fermions respectively, and $n^{b}_{i}=b^{\dagger}_{i}b_{i}(n^{f}_{i}=c^{\dagger}_{i}c_{i})$ represents the boson (fermion) number operator. $U_{bb}$ denotes the onsite repulsive interaction between bosons. We set $t_{f}=t_{b}=t=1$ as the unit of energy and $t_{ff}=3.0$ unless otherwise specified. The fermion density is fixed at $\rho_{f}=N_{f}/L=1/4$, with a fermi-boson ratio $N_{f}/N_{b}$ set to $2\colon 1$. In our density matrix renormalization group (DMRG) White (1992); Fishman et al. (2022)calculations, we consider system sizes up to $L=160$ with open boundary conditions (OBC) and $L=96$ with periodic boundary conditions(PBC). The maximum boson occupation number per site, $n^{cut}_{b}$, is set to 2 due to the boson filling $\rho_{b}=N_{b}/L=1/8$ and the large on-site repulsion $U_{bb}=8.0$. Calculations indicate that further increasing $n^{cut}_{b}$ does not alter our results.
@@
Composite trion formation.—At large values of $|t_{ff}/t|$, it has been shown Gotta et al. (2021a) that the fermions fully pair up. In such a scenario, the system exhibits only one gapless bosonic excitation mode with a central charge $c=1$. These paired fermions quasi-condensate around either $k=0$ or $k=\pi$ momenta, depending on the sign of $t_{ff}$. When bosons are introduced into the system and two-body contact interactions $U_{bf}$ are considered, the paired fermions are attracted to the bosons, leading to the formation of composite trions, as illustrated in Fig. 1. For sufficiently strong attractive interactions, $U_{bf}\ll-|t_{ff}|$, the system can be described by an effective model derived from perturbation theory Kuklov and Svistunov (2003); Mondal et al. (2021). The effective Hamiltonian is given by
$H_{eff}=-\frac{2t_{ff}t}{|U_{bf}|}\sum_{i}(c^{\dagger}_{i+1}b^{\dagger}_{i+1}c^{\dagger}_{i}c_{i}b_{i}c_{i-1}+H.c).$ (2)
@@
Correlation functions and momentum distributions.—To fully characterize the quantum phases and determine the quasi-long-range order (QLRO) within our system, we evaluate the following correlation functions: $G_{b}(r)=\langle b^{\dagger}_{i}b_{j}\rangle,G_{f}(r)=\langle c^{\dagger}_{i}c_{j}\rangle,G_{bf}(r)=\langle c^{\dagger}_{i}b^{\dagger}_{i}c_{j}b_{j}\rangle,G_{ff}(r)=\langle c^{\dagger}_{i}c^{\dagger}_{i+1}c_{j}c_{j+1}\rangle,G_{ffb}(r)=\langle c^{\dagger}_{i}b^{\dagger}_{i}c^{\dagger}_{i+1}c_{j}b_{j}c_{j+1}\rangle,$ where $r=|i-j|$ is the distance separating sites $i$ and $j$. The QLRO is determined by the power-law decay exhibited by the dominant correlation function. Notably, genuine long-range order is absent in 1D systems in the thermodynamic limit. These correlation functions take the simplest forms, either power-law decay $\sim r^{-\alpha_{q}}$ or exponential decay $\sim e^{-r/d_{q}}$, with $q$ denoting the particle species (boson $b$, fermion $f$, paired fermions $ff$, composite fermion $bf$ and composite bosonic trion $ffb$). In the context of Luttinger liquid theory, the exponents are given by $\alpha_{b(ff)}=1/(2K_{b(ff)})$, and $\alpha_{f(bf)}=(K_{f(bf)}+1/K_{f(bf)})/2$, where $K_{q}$ are the Luttinger parameters Rizzi and Imambekov (2008); Pollet et al. (2006); He et al. (2019). At zero temperature, the predominant quantum phase is determined by the correlation function with the slowest decay, characterized by the minimal positive power exponent $\alpha_{q}<2$, which signifies a divergent susceptibility Arrigoni et al. (2004); Mathey and Wang (2007).
@@
When the onsite attraction between fermions and bosons is relatively weak, the paired fermions exhibit only a weak coupling to the bosons. In this scenario, the system behaves as a two-component Luttinger liquid, comprising fermion pairs and bosons. Within this framework, the correlations of bosons, paired fermions, and composite trions all decay according to a power law $\sim r^{-\alpha}$, while the single-particle correlations of fermions and composite fermions decay exponentially, as shown in Fig. 3(a). The correlations of the trions can be understood as a direct combined projection of the correlations of the two components, arising due to the presence of $U_{bf}$ interaction. Consequently, this two-component TLL phase exhibits only two gapless bosonic modes. In accordance with (1+1)-dimensional conformal field theory (CFT) He et al. (2019); Lu et al. (2023); Calabrese and Cardy (2009, 2004); SM , this implies that the central charge should be 2. Indeed, our analysis confirms that the central charge $c\approx 2$, as determined by the von Neumann entanglement entropy at small $|U_{bf}|$.

### v4

@@ Abstract
We reveal a quantum coherent state characterized by composite bosonic trions, wherein paired fermions further bind with bosons, in one-dimensional Bose-Fermi mixtures. This phase emerges in two separate models, both featuring onsite boson–fermion attraction that induces negative binding energy for the composite trions. The first is the pair–hopping model, in which spinless fermions undergo pair hopping to form preformed pairs. The second is the extended Bose-Fermi Hubbard model, describing dipolar bosons and fermions with density–density interactions. Notably, the formation of composite trions is independent of the specific pairing interactions of fermions. Extensive density matrix renormalization group calculations demonstrate the quasi-condensation of the composite trions, evidenced by algebraically decaying correlations of trions, gapped single-particle excitations, and suppressed fermion pair correlations. Our findings provide valuable insights into the three-body pairing mechanism in mixed-particle systems.
@@ I Introduction
Three-body bound states, often referred to as trimers or trions, are fundamental entities that span diverse fields such as nuclear and particle physics, condensed matter physics, and quantum optics Hammer and Platter (2010); Guan et al. (2013); Naidon and Endo (2017). For instance, three bound quarks form neutrons and protons Peskin (2018). In few-body physics, bosonic trimers associated with the well-known Efimov effect have been extensively studied Efimov (1973); Kolganova et al. (2017); Kunitski et al. (2015). Fermionic trimers have also attracted significant interest in interacting ultracold Fermi gases Williams et al. (2009); Nakajima et al. (2011); Shi et al. (2014); Orso et al. (2010). In addition, solid-state Bose-Fermi mixtures, such as two-dimensional transition metal dichalcogenides, have emerged as promising platforms for exploring fermionic trions, where excitons (bound electron-hole pairs) bind with free electrons or holes via electrically tunable Feshbach resonances Schwartz et al. (2021); Wagner et al. (2025). Remarkably, at room temperature, Bose-Einstein condensates of exciton-polaritons have been realized in semiconductor and organic microcavities Byrnes et al. (2014); Keeling and Kéna-Cohen (2020); Ghosh et al. (2022). Polaritons are hybrid bosonic quasiparticles representing a superposition of an exciton (two fermion) and a cavity photon (one boson). Due to their finite lifetime, polaritons inherently possess a non-equilibrium nature. This naturally prompts a fundamental inquiry: in equilibrium, is there a comprehensive theoretical framework elucidating how a fermion pair and a boson can bind to form a composite bosonic trion, yielding a coherent and stable ground state? Furthermore, can such an exotic phase of matter be realized through quantum simulation?
@@
Over the past two decades, ultracold bosonic and fermionic atoms have proven to be exceptionally versatile platforms for simulating quantum physics Bloch et al. (2008); Gross and Bloch (2017); Schfer et al. (2020). Since atomic mixtures are highly controllable, rich phenomena have also been observed in Bose-Fermi mixtures (BFM), including polarons Scazza et al. (2022); Baroni et al. (2024); Fritsche et al. (2021), heteronuclear molecules Stan et al. (2004); Milczewski and Duda (2023); Duda et al. (2023), and double superfluids Ferrier-Barbut et al. (2014); Roy et al. (2017); Yao et al. (2016). Beyond extensive investigations of composite fermions (a fermion plus a boson) Stan et al. (2004); Milczewski and Duda (2023); Duda et al. (2023); Pollet et al. (2006); Cazalilla and Ho (2003); Mathey et al. (2004); Barillier-Pertuisel et al. (2008); Mering and Fleischhauer (2008); Pollet et al. (2008), previous theoretical work has analyzed the formation of fermionic trios, comprising a composite fermion and an elementary boson, under three-particle interaction using a diagrammatic formalism Kagan et al. (2004). Recently, an ultracold gas of ${}^{23}Na^{40}K_{2}$ triatomic molecules has been successfully created through an atom–diatomic-molecule Feshbach resonance between fermionic ${}^{23}Na^{40}K$ molecules and fermionic ${}^{40}K$ atoms Yang et al. (2022a, b); Cao et al. (2024). These advances in quantum simulation underscore the potential for realizing an exotic pairing process, wherein fermion pairs and bosons bind to form novel composite bosonic trions that can condense. Nevertheless, a comprehensive investigation into the underlying pairing mechanism for composite bosonic trions is still lacking.
@@
In this paper, we investigate the formation of composite bosonic trions, wherein a fermion pair further associates with a boson. We first explore the composite bosonic trion phase within the Bose–Fermi model incorporating fermion pair hopping, where the kinetic pair tunneling of spinless fermions Ruhman and Altman (2017); Gotta et al. (2021a, 2022a) induces preformed fermion pairs. Numerical evidence reveals that, under onsite boson–fermion attraction, only the trion correlation function displays algebraic decay, indicative of a single gapless bosonic mode with a central charge $c=1$. The quasi-condensation momentum of the composite trions is dictated by the sign of the fermionic pair-hopping amplitude. Furthermore, the stability of the composite trions is reinforced by negative binding energy, gapped single-particle excitations, and suppressed correlations among paired fermions. Our calculations also show that, as the boson–fermion interaction weakens, the composite trion phase transitions smoothly into a two-component Luttinger liquid (TLL), where fermion pairs and bosons each contribute a gapless excitation mode, resulting in an overall central charge of $c=2$.
@@
Building on insights gained from the preceding pair–hopping model, we generalize our analysis to the extended Bose-Fermi Hubbard model Baranov et al. (2012); Wang et al. (2010); Orignac et al. (2010); Gómez-Lozada et al. (2025), which describes dipolar bosons and fermions interacting via both short-range onsite attractions and long-range density–density interactions (including first- and second–neighbor couplings). Within this framework, we identify a similar composite trion phase that occupies a substantial portion of the phase diagram. Remarkably, the composite trion formation does not depend on the particular physical origin underlying fermion pairing, emphasizing the universal nature of the trion formation mechanism. Our work establishes a versatile platform for exploring intriguing three-body bound states, which complement the well-known bosonic trimers in Bose systems and fermionic trimers in Fermi systems, and hold promise for diverse applications in quantum many-body physics and quantum chemistry Carr et al. (2009); Bohn et al. (2017).
II Composite bosonic trions in the pair–hopping model
@@
The first model we consider incorporates the typical onsite boson-fermion interaction of strength $U_{bf}$ Pollet et al. (2006); Cazalilla and Ho (2003); Mathey et al. (2004); Barillier-Pertuisel et al. (2008); Mering and Fleischhauer (2008); Pollet et al. (2008) and a kinetic component for fermions that exhibits competition between single and paired configurations, as introduced in Ref. Gotta et al. (2021a, b, 2022b); Bariev (1991). The Hamiltonian in a chain of size $L$ is given by
$\begin{split}H=&-\sum_{i}(t_{f}c^{\dagger}_{i}c_{i+1}+t_{b}b^{\dagger}_{i}b_{i+1}+H.c.)+U_{bf}\sum_{i}n^{b}_{i}n^{f}_{i}\\ &-t_{ff}\sum_{i}(c^{\dagger}_{i+1}c^{\dagger}_{i}c_{i}c_{i-1}+H.c.)+\frac{U_{bb}}{2}\sum_{i}n^{b}_{i}(n^{b}_{i}-1),\end{split}$ (1)
@@
where $c^{\dagger}_{i}$ ($b^{\dagger}_{i}$) creates a spinless fermion (a boson) at site $i$, $t$ and $t_{ff}$ denote the tunneling rates for single particles and paired fermions respectively, and $n^{b}_{i}=b^{\dagger}_{i}b_{i}(n^{f}_{i}=c^{\dagger}_{i}c_{i})$ represents the boson (fermion) number operator. $U_{bb}$ denotes the onsite repulsive interaction between bosons. We set $t_{f}=t_{b}=t=1$ as the unit of energy and $t_{ff}=3.0$ unless otherwise specified. The fermion density is fixed at $\rho_{f}=N_{f}/L=1/4$, with a fermi-boson ratio $N_{f}/N_{b}$ set to $2\colon 1$. In our density matrix renormalization group (DMRG) White (1992); Fishman et al. (2022)calculations, we consider system sizes up to $L=160$ with open boundary conditions (OBC) and $L=96$ with periodic boundary conditions(PBC). The maximum boson occupation number per site, $n^{cut}_{b}$, is set to 2 due to the boson filling $\rho_{b}=N_{b}/L=1/8$ and the large onsite repulsion $U_{bb}=8.0$. Calculations indicate that further increasing $n^{cut}_{b}$ does not alter our results.
@@
II.1 Composite bosonic trion formation
At large values of $|t_{ff}/t_{f}|$, it has been shown Gotta et al. (2021a) that the fermions fully pair up. In such a scenario, the system exhibits only one gapless bosonic excitation mode with a central charge $c=1$. These paired fermions quasi-condensate around either $k=0$ or $k=\pi$ momenta, depending on the sign of $t_{ff}$. When bosons are introduced into the system and two-body contact interactions $U_{bf}$ are considered, the paired fermions are attracted to the bosons, leading to the formation of composite trions, as illustrated in Fig. 1. For sufficiently strong attractive interactions, $U_{bf}\ll-|t_{ff}|$, the system can be described by an effective model derived from perturbation theory Kuklov and Svistunov (2003); Mondal et al. (2021). The effective Hamiltonian is given by
$H_{eff}=-\frac{2t_{ff}t_{b}}{|U_{bf}|}\sum_{i}(c^{\dagger}_{i+1}b^{\dagger}_{i+1}c^{\dagger}_{i}c_{i}b_{i}c_{i-1}+H.c).$ (2)
@@ II.1 Composite bosonic trion formation
II.2 Correlation functions and momentum distributions
To fully characterize the quantum phases and determine the quasi-long-range order (QLRO) within our system, we evaluate the following correlation functions:
$\begin{split}&G_{b}(r)=\langle b^{\dagger}_{i}b_{j}\rangle,G_{f}(r)=\langle c^{\dagger}_{i}c_{j}\rangle,\\ &G_{bf}(r)=\langle c^{\dagger}_{i}b^{\dagger}_{i}c_{j}b_{j}\rangle,G_{ff}(r)=\langle c^{\dagger}_{i}c^{\dagger}_{i+1}c_{j}c_{j+1}\rangle,\\ &G_{ffb}(r)=\langle c^{\dagger}_{i}b^{\dagger}_{i}c^{\dagger}_{i+1}c_{j}b_{j}c_{j+1}\rangle,\end{split}$ (4)
where $r=|i-j|$ is the distance separating sites $i$ and $j$. The QLRO is determined by the power-law decay exhibited by the dominant correlation function. Notably, genuine long-range order is absent in 1D systems in the thermodynamic limit. These correlation functions take the simplest forms, either power-law decay $\sim r^{-\alpha_{q}}$ or exponential decay $\sim e^{-r/d_{q}}$, with $q$ denoting the particle species (boson $b$, fermion $f$, paired fermions $ff$, composite fermion $bf$ and composite bosonic trion $ffb$). In the context of Luttinger liquid theory, the exponents are given by $\alpha_{b(ff)}=1/(2K_{b(ff)})$, and $\alpha_{f(bf)}=(K_{f(bf)}+1/K_{f(bf)})/2$, where $K_{q}$ are the Luttinger parameters Rizzi and Imambekov (2008); Pollet et al. (2006); He et al. (2019). At zero temperature, the predominant quantum phase is determined by the correlation function with the slowest decay, characterized by the minimal positive power exponent $\alpha_{q}<2$, which signifies a divergent susceptibility Arrigoni et al. (2004); Mathey and Wang (2007a).
@@ II.2 Correlation functions and momentum distributions
When the onsite attraction between fermions and bosons is relatively weak, the paired fermions exhibit only a weak coupling to the bosons. In this scenario, the system behaves as a two-component Luttinger liquid, comprising fermion pairs and bosons. Within this framework, the correlations of bosons, paired fermions, and composite trions all decay according to a power law $\sim r^{-\alpha}$, while the single-particle correlations of fermions and composite fermions decay exponentially, as shown in Fig. 3(a). The correlations of the trions can be understood as a direct combined projection of the correlations of the two components, arising due to the presence of $U_{bf}$ interaction. Consequently, this TLL phase exhibits only two gapless bosonic modes. In accordance with (1+1)-dimensional conformal field theory (CFT) He et al. (2019); Lu et al. (2023); Calabrese and Cardy (2009, 2004), this implies that the central charge should be 2. Indeed, our analysis confirms that the central charge $c\approx 2$, as determined by the von Neumann entanglement entropy at small $|U_{bf}|$ (see Appendix B).
@@
The Fourier transforms of the correlation functions yield the momentum distributions $n_{q}(k)=\frac{1}{L}\sum_{i,j}G_{q}(i,j)e^{ik(i-j)}$ for different particles $q$. In the TLL phase, paired fermions exhibit quasi-condensation around momentum $k=0$ at $t_{ff}=-3.0$ (alternatively, around $k=\pi$ at $t_{ff}=+3.0$, as shown in Appendix B). Similarly, the momentum distribution of bosons displays a sharp peak $k=0$, as depicted in Fig. 4(a). On the contrary, in the composite trion phase, the momentum distributions of both paired fermions and bosons are significantly broadened. Notably, only the composite trions exhibit a prominent peak in momentum space, which is visible in Figure 4(b). The condensation momentum of composite trions is $\pi$ for positive $t_{ff}$ and 00 for negative $t_{ff}$, depending on the primary distribution of the constituent fermion pairs.
Figure 5: Correlation functions with $U_{bf}=-3.0$ fixed for (a) the TLL phase at $t_{ff}=-0.4$ and (b) the composite trion phase at $t_{ff}=-1.8$ in a 1D system of length $L=160$ with OBC. $\rho_{f}=2\rho_{b}=1/4$ are fixed, and $r$ represents the distance from the reference site at $L/4$. A double-logarithmic scale is used in both (a) and (b). The solid lines denote power-law fitting to $|G(r)|\sim r^{-\alpha_{q}}$ where $\alpha_{b}\simeq 0.70,\alpha_{f}\simeq 1.41,\alpha_{bf}\simeq 1.42,\alpha_{ff}\simeq 1.78,\alpha_{ffb}\simeq 1.26$ in (a) and $\alpha_{ffb}\simeq 0.67$ in (b).
II.3 Influence of magnitude of $t_{ff}$
At small $|t_{ff}/t_{f}|$, another type of two-component Luttinger liquid phase involving bosons and unpaired fermions, referred to as TLLBF, emerges. In this phase, the two gapless modes arising from the bosonic and unpaired fermionic components yield a central charge $c=2$. All relevant correlation functions—$G_{f},G_{b},G_{bf},G_{ff}$ and $G_{ffb}$—exhibit power-law decay (see Fig. 5). This behavior contrasts with that in the TLL phase at small $|U_{bf}|$ and large $|t_{ff}|$, where $G_{f}$ and $G_{bf}$ decay exponentially, while $G_{ff}$, $G_{b}$, and $G_{ffb}$ exhibit algebraic decay. A qualitative sketch of the ground-state phase diagram as a function of $t_{ff}$ and $U_{bf}$ is provided in Appendix B.
When $|t_{ff}/t_{f}|>1.6$ and the attractive $U_{bf}$ is not negligible, the composite trion phase emerges, with only composite bosonic trions displaying QLRO. This indicates the importance of preformed fermion pairs in the formation of composite bosonic trions, despite the absence of QLRO for paired fermions within the composite trion phase.